I have to use two different types of milk (eg. Milk A & B)which have two different protein to fat ratios - in percents per 100 litre of milk.
In milk A, the protein to fat ratio is 3.3/ 3.85. = .857
In milk B, the protein to fat ratio is 4.3/ 2.50 = 1.72
What quantities of milk A and B should be used to get 6000 L of mix at a ratio of (protein/fat) = 1.05
If possible let me know the equation used for this
thanks

Answer by scott8148(5895)

let A=liters of A, and B=liters of B

A+B = 6000, so A = 6000-B

A(.857) + B(1.72) = 6000(1.05)

substituting from 1st eqn. to 2nd eqn. __ (6000-B)(.857) + B(1.72)=6000(1.05)

5142 - .857B + 1.72B = 6300

subtracting 5142 __ .863B = 1158

dividing by .863 __ B = 1342L (approx)

A = 6000 - 1342 = 4658L (approx)